# MAPo : Motion-Aware Partitioning of Deformable 3D Gaussian Splatting for High-Fidelity Dynamic Scene Reconstruction

### v1

@@ 4 Method
Temporal Partitioning Based on Dynamic Scores
@@
Building on this insight, we recursively partition the 3DGs along the temporal dimension, progressively subdividing those with higher dynamic scores into finer temporal segments to achieve more precise dynamic reconstruction. For the complete time range $[0,T]$, each 3DG maintains two key properties: its partition level $l$ (initialized to 0) and its temporal segment range $t_{\text{start}},t_{\text{end}}]$ (initially $[0,T]$). For notational convenience, all such ranges are treated as left-closed and right-open. Let $G_{[t_{\text{start}},t_{\text{end}}]}$ denote the set of 3DGs assigned to the temporal segment range $t_{\text{start}},t_{\text{end}}]$. Let $F_{[t_{\text{start}},t_{\text{end}}]}$ represent the deformation network corresponding to the temporal segment range $t_{\text{start}},t_{\text{end}}]$. It is responsible for deforming the 3DGs in $G_{[t_{\text{start}},t_{\text{end}}]}$. When a 3DG at level $l$ exhibits a dynamic score exceeding the current-level threshold $\tau_{l}$ within its time range $t_{\text{start}},t_{\text{end}}]$, we partition it at the temporal midpoint $t_{\text{mid}}=(t_{\text{start}}+t_{\text{end}})/2$. The original 3DG retains the first sub-segment $t_{\text{start}},t_{\text{mid}}]$ while advancing to level $l+1$, and a new replica is created for the second sub-segment $t_{\text{mid}},t_{\text{end}}]$ with identical attributes. Correspondingly, the deformation network $F_{[t_{\text{start}},t_{\text{end}}]}$ is replicated to create $F_{[t_{\text{start}},t_{\text{mid}}]}$ and $F_{[t_{\text{mid}},t_{\text{end}}]}$ to model the distinct spatio-temporal deformation patterns within each sub-segment. Within each new shorter temporal segment, this partitioning process is applied recursively. Driven by the complexity of motion in different regions, our hierarchical approach adaptively refines their corresponding temporal segments.
Static 3D Gaussian Partitioning
@@
Cross-Frame Consistency Loss
Temporal partitioning of 3DGs, while beneficial for modeling complex motions, can introduce visual discontinuities at the partition boundaries. To ensure temporal smoothness, we introduce the cross-frame consistency loss ${L}_{\text{cross}}$, which consists of two components: ${L}_{\text{current}}$ and ${L}_{\text{gt}}$. The ${L}_{\text{current}}$ evaluates rendering consistency at the partition boundary. It compares two renderings of the same frame at this boundary:
@@
E-D3DGS 30.79 0.934 0.051 73 MB 2 hours 41 mins 37.51
Ours 31.26 0.943 0.044 65 MB 1 hour 52 mins 75.64
@@
4DGaussians 26.16 0.894 0.081 51 MB 14 mins 31.98
Ex4DGS 26.46 0.895 0.083 123 MB 1 hour 6 mins 117.49
@@
Ours 26.72 0.903 0.066 49 MB 1 hour 19 mins 92.21
Table 2: Quantitative comparison on MeetRoom dataset. Storage, training time, and FPS are calculated on discussion.
@@ 5 Experiment
Dataset and Metrics
We evaluate our method on two real-world dynamic scene datasets: N3DV (Li et al. 2022b) and Meet Room (Li et al. 2022a). The N3DV dataset includes videos (typically 10s, 30 FPS) captured by 20 cameras. For fair comparison, we downsample its images to 1352×1014 and segment the longer flame_salmon sequence into four 10s clips. The Meet Room dataset consists of 10s, 30 FPS videos (1280×720) from 13 cameras. We report PSNR, SSIM, and LPIPS for rendering quality and also assess computational costs, including training time, rendering speed, and storage, which is calculated on NVIDIA RTX A6000.
Implementation Details
Our implementation builds upon the E-D3DGS codebase. We provide comprehensive technical details in the appendix, which covers specific implementation details, hyperparameter settings, in-depth analyses of our method’s mechanics, and extensive supplementary experiments, including more baseline comparisons and additional ablation studies.
Comparisons
Quantitative Comparisons.
We compare MAPo against SOTA NeRF-based and 3DGS-based methods. To ensure fairness, all 3DGS-based baselines are evaluated under consistent conditions, including identical point cloud initializations. Direct comparisons with methods employing significantly different protocols (e.g., SWinGS, ST-GS) are deferred to the appendix, where full details on all baselines and experimental setups are also provided. As shown in Tab. 1 and Tab. 2, our method consistently achieves SOTA rendering quality across both datasets. We attribute this superior performance to our dynamic score-based partitioning strategy. Unlike approaches that rely on a single, unified deformation model (e.g., D3DGS, E-D3DGS), MAPo’s hierarchical partitioning enables specialized, independent modeling for high-dynamic 3DGs. This effectively mitigates the “temporal averaging” effect, allowing for the reconstruction of fine-grained details that are often blurred or lost in other methods, as evidenced by our top-ranking PSNR, SSIM, and LPIPS scores on both datasets. Furthermore, MAPo achieves this SOTA rendering quality while maintaining competitive computational efficiency. Our static 3D Gaussian partitioning significantly reduces overhead by eliminating redundant computations for low-dynamic 3DGs. As a result, MAPo offers a compelling balance between high fidelity and practical resource usage.
Qualitative Comparisons.
We present qualitative results in Fig. 5. The comparison highlights that baseline methods often produce degraded results in areas with complex motion. For example, in cases with fast-moving hands or detailed facial expressions, baseline methods exhibit severe motion blur and loss of detail. This is largely because their single, unified model struggles to accurately capture the diverse, complex motions. Benefiting from our temporal partitioning strategy, our approach preserves fine details in highly dynamic regions, yielding renderings that are significantly sharper and more faithful to the ground truth.
@@
Ablation Study and Analysis
We present the quantitative results of progressively incorporating each component of our method in Tab. 3. These components include: 1. Temporal Partition, referring to Temporal Partitioning Based on Dynamic Scores, where we initially employ Maximum Displacement (1.1 + Max Dis) for dynamic score computation to partition dynamic 3DGs, followed by incorporating Variance (1.2 + Var) for the full dynamic score; 2. Static Partition, denoting Static 3D Gaussian Partitioning; and 3. Cross, representing Cross-Frame Consistency Loss, which involves the sequential integration of ${L}_{\text{current}}$ (3.1 + ${L}_{\text{current}}$) and ${L}_{\text{gt}}$ (3.2 + ${L}_{\text{gt}}$). Each component is added incrementally, building upon the previously incorporated modules to evaluate their collective impact on performance. For clarity, in the subsequent experiments, we use consistent numbering to denote the variants at each stage.
@@
Figure 8: Observation of ${L}_{\text{cross}}$’s Effect on Jitter. The four images on the right are obtained by vertically concatenating the horizontal line in the “Annotation” over time.
(2.0) + Static
(3.1) + ${L}_{\text{current}}$
(3.2) + ${L}_{\text{gt}}$
Ground Truth
Figure 9: Observation of ${L}_{\text{cross}}$’s Effect on Quality.
Temporal Partition. Tab. 3 and Fig. 6 clearly demonstrate the quality improvements achieved through temporal partition. By leveraging maximum displacement to compute dynamic scores for partitioning, we significantly enhance rendering quality. Further incorporating variance into the dynamic score computation yields better rendering quality.
Static Partition. As shown in Tab. 3, Static Partition reduces computational costs compared with Temporal Partition. Additionally, Fig. 7 shows that the visual quality of static regions in Static Partition remains comparable to that of Temporal Partition without static partition.
Cross. Temporal partitioning can lead to discontinuities between frames near partition boundaries, as illustrated in Fig. 8 (2.0), where the region highlighted by a red box reveals a sharp visual discontinuity that manifests as a horizontal seam. By introducing ${L}_{\text{current}}$ and ${L}_{\text{gt}}$, we mitigate this issue. Fig. 9 shows that our cross-frame consistency loss also enhances the rendering quality, reconstructing finer details like the spray gun’s barrel, which was absent in (2.0).
@@ 6 Conclusion
We propose a novel method for high-fidelity dynamic scene reconstruction. Based on deformation-based 3DGS, MAPo leverages a dynamic score computed from historical positions to guide a hierarchical partitioning strategy that addresses the limitations of unified models. In addition, we introduce a cross-frame consistency loss to effectively mitigate visual discontinuities introduced at partition boundaries. This adaptive partitioning approach enables specialized, high-fidelity reconstruction for complex motion regions while efficiently handling static parts of the scene. Extensive experiments demonstrate that MAPo achieves state-of-the-art rendering quality, particularly excelling at capturing intricate motion details in challenging dynamic scenes while maintaining comparable computational efficiency.

### v2

@@ 4.1.1 Dynamic Score Calculation
4.1.2 Temporal Partitioning Based on Dynamic Scores
@@
Building on this insight, we recursively partition the 3DGs along the temporal dimension, progressively subdividing those with higher dynamic scores into finer temporal segments to achieve more precise dynamic reconstruction. Specifically, for the complete time range $[0,T]$, each 3DG maintains two key properties: its partition level $l$ (initialized to 0) and its temporal segment range $t_{\text{start}},t_{\text{end}}]$ (initially $[0,T]$). For notational convenience, all such ranges are treated as left-closed and right-open. Let $G_{[t_{\text{start}},t_{\text{end}}]}$ denote the set of 3DGs assigned to the temporal segment range $t_{\text{start}},t_{\text{end}}]$. Let $F_{[t_{\text{start}},t_{\text{end}}]}$ represent the deformation network corresponding to the temporal segment range $t_{\text{start}},t_{\text{end}}]$. It is responsible for deforming the 3DGs in $G_{[t_{\text{start}},t_{\text{end}}]}$. When a 3DG at level $l$ exhibits a dynamic score exceeding the current-level threshold $\tau_{l}$ within its time range $t_{\text{start}},t_{\text{end}}]$, we partition it at the temporal midpoint $t_{\text{mid}}=(t_{\text{start}}+t_{\text{end}})/2$. The original 3DG retains the first sub-segment $t_{\text{start}},t_{\text{mid}}]$ while advancing to level $l+1$, and a new replica is created for the second sub-segment $t_{\text{mid}},t_{\text{end}}]$ with identical attributes. Correspondingly, the deformation network $F_{[t_{\text{start}},t_{\text{end}}]}$ is replicated to create $F_{[t_{\text{start}},t_{\text{mid}}]}$ and $F_{[t_{\text{mid}},t_{\text{end}}]}$ to model the distinct spatio-temporal deformation patterns within each sub-segment. Within each new sub-segment, this partitioning process is applied recursively.
4.1.3 Static 3D Gaussian Partitioning
@@
4.2 Cross-Frame Consistency Loss
Temporal partitioning, while beneficial for modeling complex motions, can introduce visual discontinuities at the partition boundaries. To ensure temporal smoothness, we introduce the cross-frame consistency loss ${L}_{\text{cross}}$, which consists of two components: ${L}_{\text{current}}$ and ${L}_{\text{gt}}$. The ${L}_{\text{current}}$ evaluates rendering consistency at the partition boundary. It compares two renderings of the same frame at this boundary:
@@ 4.2 Cross-Frame Consistency Loss
LocalDyGS 30.75 0.933 0.053 102 MB 42 mins 109.30
[…]
Table 2: Quantitative comparison on the Meet Room dataset. Storage, training time, and FPS are calculated on discussion.
@@
4DGaussians 26.16 0.894 0.081 51 MB 1 hour 3 mins 77.26
[…]
@@
[…]
@@ 5 Experiment
5.1 Dataset and Metrics
We evaluate our method on two real-world dynamic scene datasets: N3DV [DyNeRF] and Meet Room [StreamRF]. The N3DV dataset includes videos at 30 FPS captured by 20 cameras. Following previous work [ed3dgs], we downsample its images to 1352×1014 and segment the longer flame_salmon sequence into four 10s clips. The Meet Room dataset consists of videos at a resolution of 1280×720 and 30 FPS from 13 cameras. We report PSNR, SSIM, and LPIPS for rendering quality, alongside computational costs including training time, rendering speed, and storage, which are calculated on an NVIDIA RTX A6000 unless otherwise specified.
5.2 Implementation Details
Our implementation builds upon the E-D3DGS codebase. In our experiments, we set the number of recorded historical positions, $m$, to 300 and the maximum partition level to 3. We provide complete implementation details in the appendix, including training specifics and more hyperparameter settings. Furthermore, we include in-depth analyses of our method’s mechanics and extensive supplementary experiments in the appendix, covering more baseline comparisons and additional ablation studies.
5.3 Comparisons
5.3.1 Quantitative Comparisons.
We compare MAPo against current open-source SOTA methods. To ensure fairness, all 3DGS-based baselines are evaluated with identical point cloud initializations. In addition to these SOTA baselines, we additionally introduce a simple segmentation baseline, E-D3DGS (seg), for comparison to highlight the advantages of our approach. This baseline applies a naive temporal partitioning strategy by uniformly splitting the video sequence into three independent segments and training a separate E-D3DGS model for each. We defer direct comparisons with methods that lack public codebases or employ significantly different protocols (e.g., SWinGS, ST-GS) to the appendix, where full details on all baselines and experimental setups are provided. As shown in Tab. 1 and Tab. 2, our method consistently achieves SOTA rendering quality across both datasets while avoiding prohibitive computational overhead, thus offering a compelling balance between high fidelity and practical resource usage.
5.3.2 Qualitative Comparisons.
We present qualitative results in Fig. 5. The comparison highlights that baseline methods often produce degraded results in areas with complex or rapid motion. For example, in cases with fast-moving hands or detailed facial expressions, baseline methods exhibit severe motion blur and loss of detail. Benefiting from our temporal partitioning strategy, our approach preserves fine details in highly dynamic regions, yielding renderings that are significantly sharper and more faithful to the ground truth.
Table 3: Progressive component ablation on Meet Room. Storage, training time, and FPS are calculated on discussion.
@@ 5.3 Comparisons
5.4 Ablation Study and Analysis
To evaluate our method, we present a progressive ablation study in Tab. 3. We establish two reference points: the E-D3DGS model, which serves as our Baseline, and a naive temporal slicing approach of E-D3DGS (Baseline (seg)) for comparison. Building upon the baseline, we incrementally add our proposed components to validate their effect. These components include: 1. Temporal Partition, our method for Temporal Partitioning Based on Dynamic Scores, where we use Maximum Displacement (1.1 + Max Dis) for dynamic score computation, then incorporate Variance (1.2 + Var); 2. Static Partition, for Static 3D Gaussian Partitioning; and 3. Cross, the Cross-Frame Consistency Loss, involving sequential integration of ${L}_{\text{current}}$ (3.1 + ${L}_{\text{current}}$) and ${L}_{\text{gt}}$ (3.2 + ${L}_{\text{gt}}$). Components are added incrementally to evaluate their collective impact. For clarity, we use consistent numbering to denote the variants at each stage.
@@ 5.4 Ablation Study and Analysis
Figure 8: Observation of ${L}_{\text{cross}}$’s effect on visual discontinuity. The four images on the right are obtained by vertically concatenating the horizontal line in the “Annotation” over time.
Figure 9: Ablation study on the $L_{cross}$. We visualize how $L_{cross}$ improves temporal consistency and rendering quality across a partition boundary (frames 74-75). The first four columns show the frame sequence to evaluate smoothness, while the fifth column provides a magnified view to compare fine detail reconstruction on the fast-moving spray gun.
Temporal Partition.
Our temporal partition strategy performs specialized temporal modeling for highly dynamic 3D Gaussians to reconstruct complex motion details that are averaged out in a single, unified model. Tab. 3 and Fig. 6 clearly demonstrate the quality improvements achieved through temporal partition. Notably, even our intermediate partitioning variants (”1.1” and ”1.2”) achieve significantly superior results compared to the naive temporal slicing baseline, Baseline (seg). This highlights that our dynamic score-based approach is far more effective at improving quality than simply training independent models on fixed temporal segments.
Static Partition.
Our static partitioning strategy leverages the dynamic score to identify and convert static 3D Gaussians, aiming to improve training efficiency, rendering speed, and storage. As shown in Tab.3, static partitioning significantly reduces computational costs compared to using temporal partitioning alone. Furthermore, Tab.3 and Fig. 7 demonstrate that applying the static partitioning strategy does not cause a decline in reconstruction quality.
Cross-Frame Consistency Loss.
We introduce the $L_{cross}$ to enforce continuity at boundaries and further enhance overall rendering quality. We quantitatively evaluate temporal consistency using the tOF [Chu_2020] metric, where lower values indicate better temporal consistency. As shown in Tab. 3, we measure tOF over both the entire video (Avg) and at the partition boundaries (Bnd). The data reveals that applying only temporal partitioning, while reducing the average tOF, causes the boundary tOF to rise significantly, which confirms the existence of visual discontinuity at the boundaries. Through the step-by-step integration of $L_{current}$ and $L_{gt}$, the boundary tOF is progressively suppressed, ultimately dropping below the baseline level in our full model and completely mitigating the issue. Our qualitative results further corroborate the dual efficacy of $L_{cross}$. As shown in Fig. 8, the time-slice visualization intuitively reveals how our method significantly mitigates the severe visual discontinuity at the boundaries. Concurrently, the consecutive frame sequence in Fig. 9 also demonstrates that our full model (+ $L_{gt}$) not only achieves the smoothest transition but also yields additional quality gains.
Analysis of Maximum Partition Level.
We conduct an ablation study on the maximum partition level to analyze its impact on reconstruction quality and computational cost. As shown in Tab. 4, progressively increasing the partition level from 0 to 5 yields a general trend of improved rendering quality. However, we observe diminishing returns in quality gains after level 3, while computational and storage costs continue to increase steadily. Therefore, we select level 3 as the maximum partition level for our main experiments to strike an optimal balance. Furthermore, the results highlight that the cost increase is manageable and not explosive; for instance, the storage at level 4 is merely double that of level 0, demonstrating the efficiency of our scheme.
Table 4: Ablation study on the partition level parameter. All experiments are conducted on the flame_salmon_frag3.
[…]
0 29.93 0.923 0.61 44MB 1 hours 13 mins 95.21
1 30.08 0.927 0.56 51MB 1 hours 24 mins 88.13
2 30.21 0.932 0.54 59MB 1 hours 37 mins 82.81
3 30.30 0.934 0.52 70MB 1 hours 56 mins 74.58
4 30.32 0.936 0.50 88MB 2 hours 22 mins 64.25
5 30.36 0.936 0.49 103MB 2 hours 40 mins 57.05
@@ 6 Conclusion
We proposed MAPo, a novel framework for high-fidelity dynamic scene reconstruction. MAPo employs a hierarchical partitioning strategy guided by a dynamic score, enabling specialized modeling for complex motion regions while treating static ones efficiently. To ensure temporal smoothness, we introduce a cross-frame consistency loss to mitigate visual discontinuities at the partition boundaries. Extensive experiments demonstrate that MAPo achieves SOTA rendering quality while maintaining competitive computational efficiency.